OpenStudy (anonymous):

Simplify i72 i –1 –i 1

OpenStudy (anonymous):

Firstly tell me when you divide 72 by 4 then what is the remainder??

OpenStudy (anonymous):

18

OpenStudy (anonymous):

I am asking for Remainder and not quotient..

OpenStudy (anonymous):

So, if Remainder = 0, then answer will be 1 Remainder = 1, then answer will be i Remainder = 2, then answer will be -1 Remainder = 3, then answer will be -i Now tell me what is the remainder when you divide 72 by 4??

OpenStudy (anonymous):

ohh i see i just figured out the problem i^72 = (i²)^36 = (-1)^36 = 1

OpenStudy (anonymous):

You can do this way also..

OpenStudy (anonymous):

I suggest you other method: \[i^4 = 1\] \[= (i^4)^{18}\] \(= 1\)

OpenStudy (anonymous):

In this case when you divide 72 by 4 you get remainder = 0, so apply the first condition I have written there..